OpenStudy (anonymous):

In how many ways can 8 instructors be assigned to six sections of a course in mathematics?

OpenStudy (anonymous):

is it possible for one instructor to teach all the sections, or can they only teach one each?

OpenStudy (anonymous):

good question!

OpenStudy (anonymous):

probably not so answer is \[8\times 7\times 6\times 5\times 4\times 3\]

OpenStudy (anonymous):

if they can teach more than one section: 8*8*8*8*8*8=262144 possibilities if they can only teach one: 8*7*6*5*4*3=20160 possibilities

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welks!